# footnotetext: $\dagger$ First authors

### v1

@@ 5 Rich and Lazy Learning
In this section we use these solutions to gain a deeper understanding of the transition between the rich and lazy regimes by examining the dynamics as a function of lambda – the relative scale - as it varies between positive and negative infinity. We investigate four key indicators of the learning regimes: the dynamics of singular vectors, the structure and robustness of the representations, and the evolution of the NTK.
Dynamics of the singular values. Here we examine a lambda-balanced linear network initialized with task-aligned weights. Previous research Saxe et al. (2019a) has demonstrated that initial weights that are aligned with the task remain aligned throughout training, restricting the learning dynamics to the singular values of the network. This setting offers a valuable opportunity to build intuition about the impact of $\lambda$ on the dynamics of learning regimes, extending beyond previous solutions Tarmoun et al. (2021); Varre et al. (2024).
@@
The proof for Theorem 5.1 can be found in Appendix C.1. As shown in Fig.3 B, as $\lambda$ approaches zero, the dynamics resemble sigmoidal learning curves that traverse between saddle points, characteristic of the rich regime Braun et al. (2022). In this regime the network learns the most salient features first, which can be beneficial for generalization Lampinen & Ganguli (2018). Conversely, as shown in Fig.3 A and C, as the magnitude of $\lambda$ increases, the dynamics become exponential, characteristic of the lazy regime. In this regime, all features are treated equally and the network’s dynamics resemble that of a shallow network. Notably, similar effects have been observed previously in the context of large absolute scales Saxe et al. (2019a) independently of the relative scale. Overall, our results highlight the critical influence relative scale $\lambda$ has in shaping the learning dynamics, from sigmoidal to exponential, steering the network between the rich and lazy regimes.
Figure 3: Simulated and analytical dynamics of the singular values of the network function with relative scale lambda A $\lambda=-2$ B $\lambda=0$ C $\lambda=2$ initialized as described in E.7.
The dynamics of the representations. We now consider how the representations of the individual parameters $\bm{W}_{1}$ and $\bm{W}_{2}$ change through training. We note that under lambda-balanced initializations there is simple structure which persists throughout training that allows us to recover the dynamics of the parameters up to a time-dependent orthogonal transformation from the dynamics of ${\bf Q}{\bf Q}^{T}(t)$.
@@
Under assumptions 4 and 4, if the network function $\bm{W}_{2}\bm{W}_{1}(t)=\bm{U}(t)\bm{S}(t)\bm{V}^{T}(t)$, then we can recover the parameters $\bm{W}_{2}(t)=\bm{U}(t)\bm{S}_{2}(t)\bm{R}^{T}(t)$ and $\bm{W}_{1}(t)=\bm{R}(t)\bm{S}_{1}(t)\bm{V}^{T}(t)$ up to time-dependent orthogonal transformation $\bm{R}(t)\in\mathbb{R}^{N_{h}\times N_{h}}$, where $\bm{S_{\lambda}}(t)=\sqrt{\bm{S}^{2}(t)+\tfrac{\lambda^{2}}{4}\mathbf{I}}$ and
@@
The effective singular values $\bm{S}_{\lambda}$ of the corresponding weights are either up-weighted or down-weighted depending on the magnitude and sign of $\lambda$, splitting the representation into two parts. This division is reflected in the network’s internal representations. With our solution, ${\bf Q}{\bf Q}^{T}(t)$, which captures the temporal dynamics of the similarity between hidden layer activations, we can analyze the network’s internal representations in relation to the task. This allows us to determine whether the network adopts a rich or lazy representation, depending on the value of $\lambda$. Assuming convergence to the global minimum, which is guaranteed when the matrix ${\bf B}$ is non-singular, the internal representation satisfies ${\bf W}_{1}^{T}{\bf W}_{1}={\bf\tilde{V}}{\bf\tilde{S}}_{1}^{2}{\bf\tilde{V}}^{T}$ and ${\bf W}_{2}{\bf W}_{2}^{T}={\bf\tilde{U}}{\bf\tilde{S}}_{2}^{2}{\bf\tilde{U}}^{T}$ with ${\bf W}_{2}{\bf W}_{1}={\bf\tilde{U}}{\bf\tilde{S}}{\bf\tilde{V}}^{T}$. Theorem C.3 in the Appendix provides a detailed proof of this limiting behavior.
@@
To illustrate this, we consider a hierarchical semantic learning task [^1] In this setting, the network has equal input and output dimensions, introduced in Saxe et al. (2014); Braun et al. (2022), where living organisms are organized according to their features (Fig. 4A). The representational similarity of the task’s inputs (${\bf\tilde{V}}{\bf\tilde{S}}{\bf\tilde{V}}^{T}$) reflects this hierarchical structure (Fig.4A). Similarly, the representational similarity of the task’s target values (${\bf\tilde{U}}{\bf\tilde{S}}{\bf\tilde{U}}^{T}$) highlights the primary groupings of items. When training a two-layer network with relative scale $\lambda$ equal to zero and task-agnostic initialization Mishkin & Matas (2015), the input and output representational similarity matrices (Fig.4 B) match the task’s structure upon convergence. As derived in Theorem C.4 the network is guaranteed to find a rich solution regardless of the absolute scale , meaning ${\bf W}_{1}^{T}{\bf W}_{1}={\bf\tilde{V}}{\bf\tilde{S}}{\bf\tilde{V}}^{T}$ and ${\bf W}_{2}{\bf W}_{2}^{T}={\bf\tilde{U}}{\bf\tilde{S}}{\bf\tilde{U}}^{T}$, as shown in Fig. 4 C. Hence the network learns task-specific representations. We also show that as $\lambda$ approaches either positive or negative infinity, the network symmetrically transitions into the lazy regime. As demonstrated in Theorem C.4 and illustrated in Fig. 4, the representations converge to an identity matrix for both large positive and large negative values of $\lambda$— emerging in the output representations for large positive $\lambda$ and input representations for large negative $\lambda$. This convergence indicates that the network adopts task-agnostic representations. Meanwhile, the other respective RSMs become negligible, with scales proportional to $1/\lambda$. Therefore, as shown in Theorem C.5, the NTK becomes static and equivalent to the identity matrix in the limit as $\lambda$ approaches infinity. However, the downscaled representations of the network remain structured and task-specific. This property could be beneficial if the weights are later rescaled, such as during fine-tuning, potentially enhancing generalization and transfer learning, as we will demonstrate in Section 6. We compare this to the scenario where both weights are initialized with large Gaussian values, leading to lazy learning that maintains a fixed NTK but lacks any structural representation, as illustrated in Fig.4. Consequently, we propose a new lazy regime, which we refer to as the semi-structured lazy regime. We note that these existing regimes preserve only the input or output representation, resulting in a partial loss of structural information. All together, we find that initialization will determine which layer in the network the task specification features resides in: layers initialized with large values will be task-agnostic, while those initialized with small values will be task-specific.
@@
Figure 4: A A semantic learning task with the SVD of the input-output correlation matrix of the task. (top) $U$ and $V$ represent the singular vectors, and $S$ contains the singular values. (bottom) The respective RSMs as $USU^{\top}$ for the input and $VSV^{\top}$ for the output task. B Simulation results and C Theoretical input and output representation matrices after training, showing convergence when initialized with values of $\lambda$ equal to $-2$, 00, and $2$, according to the initialization scheme described in E.7. D Final RSMs matrices after training converged when initialised from random large weights. E After convergence, the network’s sensitivity to input noise (top panel) is invariant to $\lambda$, but the sensitivity to parameter noise increases as $\lambda$ becomes smaller (or larger) than zero.
@@
$\left\langle\mathcal{L}\right\rangle_{\xi_{{\bf W}_{1}},\xi_{{\bf W}_{2}}}=\frac{1}{2}N_{i}\sigma^{2}_{\bf W}||{\bf W}_{2}||_{F}^{2}+\frac{1}{2}N_{o}\sigma^{2}_{\bf W}||{\bf W}_{1}||_{F}^{2}+\frac{1}{2}N_{i}N_{h}N_{o}\sigma^{4}+c,$ (19)
with norms
$||{\bf W}_{1}||_{F}^{2}=\frac{1}{2}\sum_{i=1}^{N_{h}}\left(\sqrt{4{\bf\tilde{S}}_{i}^{2}+\lambda^{2}}+\lambda\right)\quad\text{and}\quad||{\bf W}_{2}||_{F}^{2}=\frac{1}{2}\sum_{i=1}^{N_{h}}\left(\sqrt{4{\bf\tilde{S}}_{i}^{2}+\lambda^{2}}-\lambda\right).$ (20)
This implies that, under the assumption of equal input-output dimensions, networks initialized with weights such that $\lambda=0$, corresponding to the rich regime, converge to solutions that are most robust to parameter noise (Figure 4E, Appendix C.3). In practice, parameter noise could be interpreted as the noise occurring within the neurons of a biological network. Hence, a rich solution may enable a more robust representation in such systems.
The impact of the architecture. Thus far, we have found that the magnitude of the relative scale parameter $\lambda$ determines the extent or rich and lazy learning. Here, we explore how a network’s learning regime is also shaped by the interaction of its architecture and the sign of the relative scale. We consider three types of network architectures, depicted in Fig. 5A: funnel networks, which narrow from input to output ($N_{i}>N_{h}=N_{o}$); inverted-funnel networks, which expand from input to output ($N_{i}=N_{h}<N_{o}$); and square networks, where input and output dimensions are equal ($N_{i}=N_{h}=N_{o}$). Our solution, ${\bf Q}{\bf Q}^{T}$, captures the dynamics of the NTK across these different network architectures. To examine the NTK’s evolution under varying $\lambda$ initializations, we compute the kernel distance from initialization, as defined in Fort et al. (2020). As shown in Fig. 5B, we observe that funnel networks consistently enter the lazy regime as $\lambda\rightarrow\infty$, while inverted-funnel networks do so as $\lambda\rightarrow-\infty$. The NTK remains static during the initial phase, rigorously confirming the rank argument first introduced by Kunin et al. (2024) for the multi-output setting. In the opposite limits of $\lambda$, these networks transition from a lazy regime to a rich regime. During this second alignment phase, the NTK matrix undergoes changes, indicating an initial lazy phase followed by a delayed rich phase. We further investigate and quantify this delayed rich regime, showing the NTK movement over training in Fig. 5C. This behavior is also quantified in Theorem C.6, which describes the rate of learning in this network. For square networks with equal input and output dimensions, this behavior is discussed in Section 5. Across all architectures, as $\lambda\to 0$, the networks consistently transition into the rich regime. Altogether, we further characterize the delayed rich regime in wide networks.
@@ 6 Application
In this section, we apply the exact solutions for the learning dynamics in deep linear networks with lambda-balanced initialization (described in Section 4) to illustrate several specific phenomena relevant to machine learning and neuroscience.
@@
Continual learning. Continual learning, as thoroughly reviewed by Parisi et al. (2019), has long posed a significant challenge for neural network models in contrast to biological networks, particularly due to the issue of catastrophic forgetting McCloskey & Cohen (1989); Ratcliff (1990); French (1999). Similarly to the framework presented by Braun et al. (2022), our approach describes the exact solutions of the networks dynamics trained across a sequence of tasks. As detailed in Appendix D.1, we demonstrate that, regardless of the chosen value of lambda, training on subsequent tasks can result in the overwriting of previously acquired knowledge, leading to catastrophic forgetting McCloskey & Cohen (1989); Ratcliff (1990); French (1999).
@@
Reversal learning. As demonstrated in Braun et al. (2022), reversal learning theoretically does not succeed in deep linear networks as the initalization aligns with the separatrix of a saddle point. While simulations show that the learning dynamics can escape the saddle point due to numerical imprecision, the process is catastrophically slowed in its vicinity. However, when $\lambda$ is non-zero, reversal learning dynamics consistently succeed, as they avoid passing through the saddle point due to the initialization scheme. This is both theoretically proven and numerically illustrated in Appendix D.2. We also present a spectrum of reversal learning behaviors controlled by the relative scale $\lambda$, ranging from rich to lazy learning regimes. This spectrum has the potential to explain the diverse dynamics observed in animal behavior, offering insights into the learning regimes relevant to various neuroscience experiments.
@@
Transfer learning. We consider how different $\lambda$ initializations influence generalization to a new feature after being trained on an initial task. As detailed in Appendix D.3 we first train each network on the hierarchical semantic learning task described in Fig. 4. After, we add a new feature to the dataset for example ‘eats worms’ We train it specifically on the corresponding item, in this case, the goldfish, while keeping the rest of the network parameters unchanged. Afterwards, we evaluate the generalization to the other items. We observe in Appendix figure 8 that the hierarchical structure of the data is effectively transferred to the new feature when the representation is task-specific and $\lambda$ is zero. Conversely, when the output feature representation is lazy, meaning the hidden representation lacks adaptation, no hierarchical generalization is observed. Strikingly, when $\lambda$ is positive, the hierarchical structure in the input weights remains small but structured, while the output weights exhibit a lazy representation and the network generalizes hierarchically. This indicates that the lazy regime structure can be beneficial for transfer learning.

### v2

@@ 5 Rich and Lazy Learning
Next, we use these solutions to gain a deeper understanding of the transition between the rich and lazy regimes by examining the dynamics as a function of $\lambda$ – the relative scale – as it varies between positive and negative infinity. We investigate four key indicators of the learning regimes: the dynamics of singular vectors, the structure and robustness of the representations, and the evolution of the NTK.
Dynamics of the singular values. Here we examine a $\lambda$-balanced linear network initialized with task-aligned weights. Previous research (Saxe et al., 2019a) has demonstrated that initial weights that are aligned with the task remain aligned throughout training, restricting the learning dynamics to the singular values of the network. This setting offers a valuable opportunity to build intuition about the impact of $\lambda$ on the dynamics of learning regimes, extending beyond previous solutions (Tarmoun et al., 2021; Varre et al., 2024).
@@
The proof for Theorem 5.1 can be found in Appendix C.1. As shown in Fig. 3B, as $\lambda$ approaches zero, the dynamics resemble sigmoidal learning curves that traverse between saddle points, characteristic of the rich regime (Braun et al., 2022). In this regime the network learns the most salient features first, which can be beneficial for generalization (Lampinen & Ganguli, 2018). Conversely, as shown in Fig. 3A and C, as the magnitude of $\lambda$ increases, the dynamics become exponential, characteristic of the lazy regime. In this regime, all features are treated equally and the network’s dynamics resemble that of a shallow network. Notably, similar effects have been observed previously in the context of large absolute scales (Saxe et al., 2019a) independently of the relative scale. Overall, our results highlight the critical influence the relative scale $\lambda$ has in shaping the learning dynamics, from sigmoidal to exponential, steering the network between the rich and lazy regimes.
Figure 3: Simulated and analytical dynamics of the singular values of the network function with relative scale of A. $\lambda=-2$, B. $\lambda=0$, or C. $\lambda=2$, initialized as described in Appendix E.7.
The dynamics of the representations. We now consider how the representations of the individual parameters $\bm{W}_{1}$ and $\bm{W}_{2}$ change through training. We note that under $\lambda$-balanced initializations there is a simple structure which persists throughout training that allows us to recover the dynamics of the parameters up to a time-dependent orthogonal transformation from the dynamics of ${\bf Q}{\bf Q}^{T}(t)$.
@@
Under assumptions 4, if the network function $\bm{W}_{2}\bm{W}_{1}(t)=\bm{U}(t)\bm{S}(t)\bm{V}^{T}(t)$ is full rank, then we can recover the parameters $\bm{W}_{2}(t)=\bm{U}(t)\bm{S}_{2}(t)\bm{R}^{T}(t)$ and $\bm{W}_{1}(t)=\bm{R}(t)\bm{S}_{1}(t)\bm{V}^{T}(t)$ up to time-dependent orthogonal transformation $\bm{R}(t)\in\mathbb{R}^{N_{h}\times N_{h}}$, where $\bm{S_{\lambda}}(t)=\sqrt{\bm{S}^{2}(t)+\tfrac{\lambda^{2}}{4}\mathbf{I}}$ and
@@
The effective singular values $\bm{S}_{\lambda}$ of the corresponding weights are either up-weighted or down-weighted depending on the magnitude and sign of $\lambda$, splitting the representation into two parts. This division is reflected in the network’s internal representations. With our solution, ${\bf Q}{\bf Q}^{T}(t)$, which captures the temporal dynamics of the similarity between hidden layer activations, we can analyze the network’s internal representations in relation to the task. This allows us to determine whether the network adopts a rich or lazy representation, depending on the value of $\lambda$. Assuming convergence to the global minimum, which is guaranteed when the matrix ${\bf B}$ is non-singular, the internal representation satisfies ${\bf W}_{1}^{T}{\bf W}_{1}={\bf\tilde{V}}{\bf\tilde{S}}_{1}^{2}{\bf\tilde{V}}^{T}$ and ${\bf W}_{2}{\bf W}_{2}^{T}={\bf\tilde{U}}{\bf\tilde{S}}_{2}^{2}{\bf\tilde{U}}^{T}$ with ${\bf W}_{2}{\bf W}_{1}={\bf\tilde{U}}{\bf\tilde{S}}{\bf\tilde{V}}^{T}$. Theorem C.3 in the Appendix provides a detailed proof of this limiting behavior. To illustrate this, we consider a hierarchical semantic learning task [^1] In this setting, the network has equal input and output dimensions, introduced in Saxe et al. (2014); Braun et al. (2022), where living organisms are organized according to their features (Fig. 4A). The representational similarity of the task’s inputs (${\bf\tilde{V}}{\bf\tilde{S}}{\bf\tilde{V}}^{T}$) reflects this hierarchical structure (Fig.4A). Similarly, the representational similarity of the task’s target values (${\bf\tilde{U}}{\bf\tilde{S}}{\bf\tilde{U}}^{T}$) highlights the primary groupings of items. When training a two-layer network with relative scale $\lambda$ equal to zero and task-agnostic initialization (Mishkin & Matas, 2015), the input and output representational similarity matrices (Fig.4B) match the task’s structure upon convergence. As derived in Theorem C.4 the network is guaranteed to find a rich solution regardless of the absolute scale , meaning ${\bf W}_{1}^{T}{\bf W}_{1}={\bf\tilde{V}}{\bf\tilde{S}}{\bf\tilde{V}}^{T}$ and ${\bf W}_{2}{\bf W}_{2}^{T}={\bf\tilde{U}}{\bf\tilde{S}}{\bf\tilde{U}}^{T}$, as shown in Fig. 4C. Hence, the network learns task-specific representations. We also show that as $\lambda$ approaches either positive or negative infinity, the network symmetrically transitions into the lazy regime. As demonstrated in Theorem C.4 and illustrated in Fig. 4, the representations converge to an identity matrix for both large positive and large negative values of $\lambda$— emerging in the output representations for large positive $\lambda$ and input representations for large negative $\lambda$. This convergence indicates that the network adopts task-agnostic representations. Meanwhile, the other respective RSMs become negligible, with scales proportional to $1/\lambda$. Therefore, as shown in Theorem C.5, the NTK becomes static and equivalent to the identity matrix in the limit as $\lambda$ approaches infinity. However, the downscaled representations of the network remain structured and task-specific. Intuitively, in this setup, the larger weights function as an identity-like projection, while the smaller weights adapt and align to the task. However, due to their relative scale compared to the larger weights, their contribution to the NTK remains negligible. This property could be beneficial if the weights are later rescaled, such as during fine-tuning, potentially enhancing generalization and transfer learning, as we will demonstrate in Section 6. We compare this to the scenario where both weights are initialized with large Gaussian values, leading to lazy learning that maintains a fixed NTK but lacks any structural representation, as illustrated in Fig. 4D. We further discuss the relationship between the scale and the relative scale in appendix A.4. Furthermore, in the infinite-width regime, where weights are initialized from a Gaussian distribution with large variance, averaging effects cause both input and output representations to approximate identity matrices. In this scenario, the network learns in the lazy regime with a fixed Neural Tangent Kernel (NTK). This behavior contrasts with the dynamics observed in the current setting since both input and output representations are task agnostic. Consequently, we propose a new lazy regime, which we refer to as the semi-structured lazy regime. We note that these existing regimes preserve only the input or output representation, resulting in a partial loss of structural information. In the nonlinear setting, this behavior is not expected to hold, as an additional factor comes into play in the computation of the NTK: the activation coefficients of the nonlinearity, as demonstrated in Kunin et al. (2024). In that case large relative weight (large positive $\lambda$) leads to a rapid rich regime. All together, we find that initialization will determine which layer in the network the task specification features resides in: layers initialized with large values will be task-agnostic, while those initialized with small values will be task-specific.
@@
Figure 4: A. A semantic learning task with the SVD of the input-output correlation matrix of the task. (top) $U$ and $V$ represent the singular vectors, and $S$ contains the singular values. (bottom) The respective RSMs as $USU^{\top}$ for the input and $VSV^{\top}$ for the output task. B. Simulation results and C. Theoretical input and output representation matrices after training, showing convergence when initialized with values of $\lambda$ equal to $-2$, 00, and $2$, according to the initialization scheme described in Appendix E.7. D. Final RSMs matrices after training converged when initialised from random large weights. E. After convergence, the network’s sensitivity to input noise (top panel) is invariant to $\lambda$, but the sensitivity to parameter noise increases as $\lambda$ becomes smaller (or larger) than zero.
@@
The impact of the architecture. Thus far, we have found that the magnitude of the relative scale parameter $\lambda$ determines the extent of rich and lazy learning. Here, we explore how a network’s learning regime is also shaped by the interaction of its architecture and the sign of the relative scale. We consider three types of network architectures, depicted in Fig. 5A: funnel networks, which narrow from input to output ($N_{i}>N_{h}=N_{o}$); inverted-funnel networks, which expand from input to output ($N_{i}=N_{h}<N_{o}$); and square networks, where input and output dimensions are equal ($N_{i}=N_{h}=N_{o}$). Our solution, ${\bf Q}{\bf Q}^{T}$, captures the dynamics of the NTK across these different network architectures. To examine the NTK’s evolution under varying $\lambda$ initializations, we compute the kernel distance from initialization, as defined in Fort et al. (2020). As shown in Fig. 5B, we observe that funnel networks enter the lazy regime as $\lambda\rightarrow\infty$, while inverted-funnel networks do so as $\lambda\rightarrow-\infty$. The NTK remains static during the initial phase, rigorously confirming the rank argument first introduced by Kunin et al. (2024) for the multi-output setting. In the opposite limits of $\lambda$, these networks transition from a lazy regime to a rich regime. During this second alignment phase, the NTK matrix undergoes changes, indicating an initial lazy phase followed by a delayed rich phase. We further investigate and quantify this delayed rich regime, showing the NTK movement over training in Fig. 5C. This behavior is also quantified in Theorem C.6, which describes the rate of learning in this network. In Appendix A.4, we further explore the impact of the architecture as a function of the absolute and relative scale. Intuitively, the delayed onset of the rich regime occurs because no least-squares solution exists within the span of the network at initialization. In such cases, the network enters a delayed rich phase, where $\lambda$ tends toward infinity, with the magnitude of $\lambda$ determining the length of the delay. At first, the network exhibits lazy dynamics, striving to approximate the solution. However, as constraints necessitate adjustments in its directions, the network gradually transitions into the rich phase. For square networks this behavior is discussed in Section 5. Across all architectures, as $\lambda\to 0$, the networks consistently transition into the rich regime. Altogether, we further characterize the delayed rich regime in wide networks.
@@ 6 Applications
In this section, we apply the exact solutions for the learning dynamics in deep linear networks described in Section 4 to illustrate several phenomena relevant to machine learning and neuroscience.
@@
Continual learning. Continual learning, as thoroughly reviewed by Parisi et al. (2019), has long posed a significant challenge for neural network models in contrast to biological networks, particularly due to the issue of catastrophic forgetting (McCloskey & Cohen, 1989; Ratcliff, 1990; French, 1999). Similarly to the framework presented by Braun et al. (2022), our approach describes the exact solutions of the networks dynamics trained across a sequence of tasks describing the entire continual learning process. As detailed in Appendix D.1, we demonstrate that, regardless of the chosen value of $\lambda$, training on subsequent tasks can result in the overwriting of previously acquired knowledge, leading to catastrophic forgetting.
@@
Reversal learning. During reversal learning, previously acquired knowledge must be relearned, necessitating the overcoming of an earlier established relationship between inputs and outputs. As demonstrated in Braun et al. (2022), reversal learning theoretically does not succeed in deep linear networks as the initalization aligns with the separatrix of a saddle point. While simulations show that the learning dynamics can escape the saddle point due to numerical imprecision, the process is catastrophically slowed in its vicinity. However, when $\lambda$ is non-zero, reversal learning dynamics consistently succeed, as they avoid passing through the saddle point due to the initialization scheme. This is both theoretically proven and numerically illustrated in Appendix D.2. We also present a spectrum of reversal learning behaviors controlled by the relative scale $\lambda$, ranging from rich to lazy learning regimes. This spectrum has the potential to explain the diverse dynamics observed in animal behavior, offering insights into the learning regimes relevant toneuroscience experiments.
@@
Transfer learning. We consider how different $\lambda$ initializations influence generalization to a new feature after being trained on an initial task. As detailed in Appendix D.3, we first train each network on the hierarchical semantic learning task described in Fig. 4. We then add a new feature to the dataset, and train the network specifically on the corresponding item while keeping the rest of the network parameters unchanged. Afterwards, we evaluate the generalization to the other items. We observe in Appendix Fig. 13 that the hierarchical structure of the data is effectively transferred to the new feature when the representation is task-specific and $\lambda$ is zero. Conversely, when the input feature representation is lazy ($\lambda\leq 0$), meaning the hidden representation lacks adaptation, no hierarchical generalization is observed. Strikingly, when $\lambda$ is positive, the hierarchical structure in the input weights remains small but structured, while the output weights exhibit a lazy representation and the network generalizes hierarchically. Specifically Fig. 13 shows that the generalization loss on untrained items with the new feature decreases as a function of increasing $\lambda$. Therefore, as $\lambda$ increases, networks more effectively transfer the hierarchical structure of the network to the new feature for untrained items, leading to an increase in generalization performance. This indicates that the lazy regime structure (large $\lambda$ values) can be beneficial for transfer learning.
Fine-tuning It is a common practice to pre- train neural networks on a large auxiliary task before fine-tuning them on a downstream task with limited samples. Despite the widespread use of this approach, the dynamics and outcomes of this method remain poorly understood. Our study establishes a theoretical basis for the success of fine-tuning, particularly how changes in $\lambda$-balancedness initialisation after pre-training affect performance on new datasets (see Appendix D.4). We consistently find that finetuning performance improves and converges more quickly as networks are re-balanced to larger values of $\lambda_{FT}$ and, conversely, decreases as $\lambda_{FT}$ approaches 0 as shown in Fig. 14. Our work examines the fine-tuning dynamics of two-layer linear networks. While simple, these architectures are commonly utilized for fine-tuning large pre-trained language and vision models, notably in Low-Rank Adapters (LoRA) (Hu et al., 2022) as further discussed in Appendix D.4. While a detailed exploration of fine-tuning performance in practice as a function of initialization lies beyond the scope of this work, it remains an important direction for future research.